OpenStudy (anonymous):

HELP in 6th grade –7 + g = 13 A. –20 B. –6 C. 6 D. 20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D(20), if yo uput -7 on to the other other side it becomes 13+7=20

OpenStudy (anonymous):

thanks can you guys help me with more

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

u have to explain otherwise @angelv00 wont understand when the finals come up like exams no one will give @AngelRae answer so u have to explain

OpenStudy (anonymous):

Solve. k + (–7) = –3 A. –10 B. –4 C. 4 D. 10

OpenStudy (anonymous):

C

OpenStudy (anonymous):

can some one explain this please

OpenStudy (anonymous):

since u have 2 negatives u change the sign into subtraction which would give u a positive because of the 2 negatives

OpenStudy (shamil98):

you all lie it's obviously 99

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

+ and - = - so 4-7=-3

OpenStudy (anonymous):

yes perfect sense

OpenStudy (anonymous):

so is A the answer

OpenStudy (anonymous):

no C

OpenStudy (anonymous):

a negative and positive will give u a negative number

OpenStudy (anonymous):

here is another one

OpenStudy (anonymous):

the answer would be C because u have to change the sign into subtraction thanks to the 2 negatives

OpenStudy (anonymous):

–24 = –31 + w A. –55 B. –7 C. 7 D. 55

OpenStudy (anonymous):

C for the second one

OpenStudy (anonymous):

inverse the -24 too the other side which becomes +24, so -31 +24=7

OpenStudy (anonymous):

i think the answer is C

OpenStudy (anonymous):

n + (–14) = 9 A. –23 B. –5 C. 5 D. 23

OpenStudy (anonymous):

i think the answer is -5

OpenStudy (anonymous):

14+9=23 so D

OpenStudy (anonymous):

i passed my test

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@Starr_DynastyT You Tagged Me Why?

OpenStudy (anonymous):

oh didnt mean to sorry about that it was a mistake @AngelRae

OpenStudy (anonymous):

@Starr_DynastyT Its Cool Beans ^.^

OpenStudy (anonymous):

coolio @AngelRae